How many different committees can be formed from 7 teachers and 9 students if the committee is to consist of 3 teachers and 4​ students?
Answer by math_helper(2461)

You are choosing 3 of 7 teachers, that can be done in 7C3 = (7!)/(3!*4!) = 35 ways
and you are choosing 4 of 9 students, that can be done in 9C4 = (9!)/(4!*5!) = 126 ways
The total committees that can be formed is then 35*126 = 4410